# Technological Complexity Based on Japanese Patent Data

### v3

@@ Discussion
This study applies the Technological Complexity Index (TCI) to Japanese corporate patent data across various technological fields to identify sophisticated areas critical to innovation. Accurate assessment of technological complexity is essential for policymakers and corporate strategists aiming to maintain competitive advantages amidst intensifying global competition. Previous regional approaches have struggled with instability when employing detailed technological classifications, resulting in coarse and potentially misleading insights. This research provides three primary contributions. First, it demonstrates methodological robustness at the corporate level. Unlike regional approaches, corporate-level TCI retains stable technological rankings even when classifications become highly detailed. Our results reveal notable complexity in the Pharmaceuticals and Chemistry sectors, likely reflecting Japan’s industrial focus during the pharmaceutical industry’s formative period. This consistency across classification scales represents significant progress compared to previous regional analyses that lost reliability when switching from broader to finer classifications. Second, the empirical findings underscore the advantage of evaluating technological complexity at the corporate level. Prior research suggested that complexity rankings could converge due to methodological limitations of classification systems [8]. However, our corporate-level analysis maintains clear distinctions between technological fields even at granular classification levels. This evidence highlights the corporate approach’s potential to provide deeper and more accurate insights into technological dynamics within specific industries or regions. Third, our results offer practical implications for industrial policy and strategy. The identification of high-complexity technological sectors informs targeted investment and innovation strategies, especially relevant amid rapid advances in artificial intelligence and growing global competition. Corporate strategists can leverage these insights to allocate resources effectively, fostering technological leadership and competitive differentiation. Nevertheless, our study faces several limitations. Capturing dynamic shifts in technological trends requires a balance between long-term analysis, which ensures stability, and short-term assessments, which detect emerging innovations. Shortening aggregation periods introduces challenges due to the sensitivity of TCI calculations to sparse patent data, thus decreasing result stability [27]. The HH algorithm underlying TCI calculations evaluates complexity based on node connections within a bipartite network, making it vulnerable to inaccuracies when patent data are sparse. To address these challenges, future studies could integrate advanced methods such as the Fitness Complexity algorithm [13], which employs nonlinear iterative processes to better capture network structures, enhancing robustness and reliability [14, 15]. Additionally, weighting links by citation counts rather than raw patent numbers could offer a more precise representation of technological effort and impact. Future research directions include comparative robustness analyses of TCI with alternative indicators such as patent citations, application of machine learning techniques to predict technological shifts, and cross-national comparative studies. Such explorations will deepen our understanding of technological complexity and provide richer guidance for industrial development strategies. Ultimately, this research significantly advances our understanding of technological complexity at the corporate level in Japan. By clearly identifying key technology sectors, our findings inform strategic planning and policy-making, positioning Japan advantageously in an era marked by technological acceleration and intensified international competition.
@@ Methods
The dataset utilized in this study comprises registered patents submitted to the Japan Patent Office (JPO) covering fiscal years 1981 through 2010, encompassing detailed records of 3,189,536 patents filed by 64,622 corporations. Patent records adhere to the International Patent Classification (IPC) system and are further categorized according to the Schmoch classification [21].
@@
The Schmoch system aggregates IPC codes into 35 distinct technological fields, ensuring balanced class sizes and technological consistency, and has been widely applied in previous research [29, 20, 30]. Nevertheless, the granularity of classification substantially influences analytical outcomes. Administrative classifications vary markedly in granularity and international comparability, potentially introducing analytical noise or biases if inadequately managed [8]. Employing consistent classification schemes, exemplified by the Schmoch system utilized in this study, mitigates such issues by providing balanced and coherent technological categories for robust analyses.
@@
Traditional methods for measuring patent contributions typically employ raw patent counts [31, 20, 29, 32, 30, 17, 33]. However, such counting methods can lead to overestimation, particularly when patents involve multiple corporations or span multiple technological domains [29]. To address potential overestimations, the primary IPC class of each patent is mapped onto the corresponding Schmoch field, with patent value fractionally allocated by equally dividing it among associated corporations and technological fields. This fractional allocation prevents overestimation arising from multi-corporation ownership or cross-domain classifications.
@@
Additionally, filtering criteria are essential to exclude extreme cases capable of distorting analytical outcomes [8, 29]. Accordingly, a filtering criterion based on patent distribution per corporation is applied in this study, retaining only the top 3% of corporations in patent output (Fig. A3). This filtering criterion excludes corporations exhibiting minimal patenting activity, thereby reducing significant analytical noise in network analyses. Following this filtering, the dataset comprises 2,485,027 patents filed by 1,938 corporations, representing approximately 93% of the original patent volume.
@@
Utilizing these filtered data, technological complexity is based on a bipartite network connecting corporations to technological fields in which they are specialized. Although there are some arguments regarding the optimal specialization index for establishing connections within these networks [29], we follows prior research by employing the Revealed Technological Advantage (RTA) index [34]. The RTA index builds upon the revealed comparative advantage framework proposed by Balassa [35]. Specifically,The RTA index builds upon the revealed comparative advantage framework proposed by Balassa [35]. Specifically,the RTA index for a given corporation in a specific technological field measures the ratio between the share of patents held by corporation $C$ in technology $T$ and the share of patents in technology $T$ among all corporations. The numerator of the RTA index captures the fraction of patent activity devoted to technology $T$ for corporation $C$. Similarly, the denominator reflects the overall share of patent activity in technology $T$ among all corporations, computed as the sum of $W_{CT}$ over all corporations divided by the sum of $W_{CT}$ over all corporations and technologies. The RTA index is thus given by
$\mbox{RTA}_{CT}=\frac{W_{CT}}{\sum_{T}W_{CT}}\Bigg{/}\frac{\sum_{C}W_{CT}}{\sum_{CT}W_{CT}}\leavevmode\nobreak\ .$ (1)
@@
Here, the metric $K_{T,1}$ represents the average nearest neighbor degree for technology nodes and it reflects the average diversity of corporations connected to a given technology. Substituting Eq. 4 into Eq. 5 yields
@@
$\widetilde{M}_{TT^{\prime}}=\sum_{C}\frac{M_{CT}\,M_{CT^{\prime}}}{K_{C,0}\,K_{T,0}}\leavevmode\nobreak\ .$ (7)
Eq. 6 is satisfied when $K_{T,N}=K_{T,N-2}=1$, which corresponds to the eigenvector associated with the largest eigenvalue of the stochastic matrix $\widetilde{M}_{T,T^{\prime}}$. Since this eigenvector is a vector of ones and not informative, the eigenvector corresponding to the second largest eigenvalue of $\widetilde{M}_{TT^{\prime}}$ captures the greatest variance among technology fields [8, 36]. Thus, TCI is given by
@@
where $\widetilde{T}$ represents the eigenvector corresponding to the second largest eigenvalue of $\widetilde{M}_{TT^{\prime}}$, and $\langle\widetilde{T}\rangle$ is their mean and $\text{stdev}(\widetilde{T})$ is their standard deviation.
@@
All scripts and code used to compute the Technological Complexity Index and reproduce the results of this paper are publicly available in the following GitHub repository: https://github.com/RintaroKARASHIMA/KCIinJapaneseFirms. The implementation of the Economic Complexity algorithms is based on the py-ecomplexity library (MIT License), available at https://github.com/cid-harvard/py-ecomplexity.

### v5

@@ Discussion
This study calculates the Technological Complexity Index (TCI) using Japanese patent data identify sophisticated technologies critical to innovation, and provides three primary contributions. First, it demonstrates methodological robustness at the corporate level. Unlike regional approaches, corporate-level TCI retains consitent technological rankings even when classifications become highly detailed. Our results reveal notable complexity in the Pharmaceuticals and Chemistry sectors, likely reflecting Japan’s industrial focus during the pharmaceutical industry’s formative period. This consistency across classification scales represents significant progress compared to previous regional analyses that lost reliability when switching from broader to finer classifications. Second, by visualizing the temporal evolution of rankings, this study clarifies which technologies appear to have sustained innovative activity during Japan’s prolonged period of modest economic growth. The corporate level lens makes these trajectories easier to read because rankings remain stable enough to track long run patterns while still capturing meaningful changes. This perspective highlights fields that persist at high complexity as well as those that rise or fade, providing a more concrete basis for interpreting how Japan’s innovation system has been supported over time. Third, the corporate level approach also enables a finer view of regional technological characteristics. By attributing firms to places and then examining their technological portfolios, we can recover local profiles that are obscured by coarse regional aggregates. This yields clearer distinctions among prefectures and urban areas and helps identify locally salient capabilities that may be missed when analysis is performed only at the regional level.
At the same time, there are limits that point to future work. Detecting dynamic shifts requires a balance between longer windows that improve stability and shorter windows that capture emerging fields. Our TCI is computed on a bipartite network in which corporations are connected to technology classes. A link is present when a corporation is active in a class according to an operational rule, for example filing a sufficient number of patents in that class or meeting an RCA based criterion. When patent counts are sparse, the network becomes thin and the HH iterations that infer complexity from the pattern of connections can become unstable [PintarEssletzbichler2022]. To mitigate this sparsity problem, future work can refine both the link rule and the information carried by each link. One direction is to move from unweighted to weighted edges that reflect the intensity and impact of inventive activity. Citation based measures are natural candidates, such as raw citations, family normalized citations, or time decayed citations, which may represent technological effort and impact more precisely than simple counts. Another direction is to adopt alternative iterative schemes such as the Fitness Complexity algorithm [Tacchella2012], which uses non linear iterations to capture structure in sparse settings and may yield more robust rankings [Wu2016, Albeaik2017].
Building on these methodological directions, future research should examine how such refinements affect the robustness and generality of TCI. Comparative analyses can evaluate TCI against citation based indicators and alternative weighting schemes. Machine learning models can then use these enriched features to predict shifts in technological rankings. Cross national applications will test whether the corporate level stability observed for Japan also appears in other innovation systems. Ultimately, this study advances the corporate level understanding of technological complexity in Japan. By identifying high complexity technology fields with greater consistency, it provides actionable evidence for strategic planning and policy design in a period of rapid technological change and intensifying global competition.
@@ Methods
The dataset utilized in this study comprises registered patents submitted to the Japan Patent Office (JPO) covering fiscal years 1981 through 2010, encompassing detailed records of 3,189,536 patents filed by 64,622 corporations. Patent records adhere to the International Patent Classification (IPC) system and are further categorized according to the Schmoch classification [Schmoch2008].
@@
The Schmoch system aggregates IPC codes into 35 distinct technological fields, ensuring balanced class sizes and technological consistency, and has been widely applied in previous research [PintarScherngell2022, Balland2018, Whittle2019]. Nevertheless, the granularity of classification substantially influences analytical outcomes. Administrative classifications vary markedly in granularity and international comparability, potentially introducing analytical noise or biases if inadequately managed [Hidalgo2021]. Employing consistent classification schemes, exemplified by the Schmoch system utilized in this study, mitigates such issues by providing balanced and coherent technological categories for robust analyses.
@@
Traditional methods for measuring patent contributions typically employ raw patent counts [Balland2016, Balland2018, PintarScherngell2022, Pinheiro2022, Whittle2019, Jun2023, Abay2024]. However, such counting methods can lead to overestimation, particularly when patents involve multiple corporations or span multiple technological domains [PintarScherngell2022]. To address potential overestimations, the primary IPC class of each patent is mapped onto the corresponding Schmoch field, with patent value fractionally allocated by equally dividing it among associated corporations and technological fields. This fractional allocation prevents overestimation arising from multi-corporation ownership or cross-domain classifications.
The regional approach evaluates technological complexity as a relative index that reflects regional technological capabilities. However, this method is sensitive to spatial disparities and tends to yield less stable results when finer classifications such as IPC classes are used, compared to broader classifications like the Schmoch system [PintarEssletzbichler2022]. Such instability appears in both regional and technological complexity evaluations, as it is inherent in the symmetric structure of the underlying bipartite network.
@@
In this study, we analyze technological complexity at the corporate level to overcome these limitations and achieve a more detailed understanding of regional technological portfolios. Additionally, filtering criteria are essential to exclude extreme cases capable of distorting analytical outcomes [Hidalgo2021, PintarScherngell2022]. Accordingly, a filtering criterion based on patent distribution per corporation is applied in this study, retaining only the top 3% of corporations in patent output (Fig. S3). This filtering criterion excludes corporations exhibiting minimal patenting activity, thereby reducing significant analytical noise in network analyses. Following this filtering, the dataset comprises 2,900,359 patents filed by 1,939 corporations, representing approximately 91% of the original patent volume.
@@
Utilizing these filtered data, technological complexity is based on a bipartite network connecting corporations to technological fields in which they are specialized. Although there are some arguments regarding the optimal specialization index for establishing connections within these networks [PintarScherngell2022], we follows prior research by employing the Revealed Technological Advantage (RTA) index [Soete1987]. The RTA index builds upon the revealed comparative advantage framework proposed by Balassa [Balassa1965]. Specifically, the RTA index builds upon the revealed comparative advantage framework proposed by Balassa [Balassa1965]. Specifically,the RTA index for a given corporation in a specific technological field measures the ratio between the share of patents held by corporation $C$ in technology $T$ and the share of patents in technology $T$ among all corporations. The numerator of the RTA index captures the fraction of patent activity devoted to technology $T$ for corporation $C$. Similarly, the denominator reflects the overall share of patent activity in technology $T$ among all corporations, computed as the sum of $W_{CT}$ over all corporations divided by the sum of $W_{CT}$ over all corporations and technologies. The RTA index is thus given by
$\mbox{RTA}_{CT}=\frac{W_{CT}}{\sum_{T}W_{CT}}\Bigg/\frac{\sum_{C}W_{CT}}{\sum_{CT}W_{CT}}\penalty 10000\ .$ (1)
@@
Here, the metric $K_{T,1}$ represents the average nearest neighbor degree for technology nodes and it reflects the average diversity of corporations connected to a given technology. Substituting Eq. (4) into Eq. (5) yields
@@
$\widetilde{M}_{TT^{\prime}}=\sum_{C}\frac{M_{CT}\,M_{CT^{\prime}}}{K_{C,0}\,K_{T,0}}\penalty 10000\ .$ (7)
Eq. (6) is satisfied when $K_{T,N}=K_{T,N-2}=1$, which corresponds to the eigenvector associated with the largest eigenvalue of the stochastic matrix $\widetilde{M}_{TT^{\prime}}$. Since this eigenvector is a vector of ones and not informative, the eigenvector corresponding to the second largest eigenvalue of $\widetilde{M}_{TT^{\prime}}$ captures the greatest variance among technology fields [Hidalgo2021, Mealy2019]. Thus, TCI is given by
@@
[…]
@@
All scripts and code used to compute the Technological Complexity Index and reproduce the results of this paper are publicly available in the following GitHub repository: https://github.com/RintaroKARASHIMA/KCIinJapaneseFirms.
Author Contributions Statement
H.I. conceived the main research idea, designed the study, collected the data, and secured the funding. R.K. performed the analysis and drafted the manuscript. Both H.I. and R.K. critically reviewed the results, discussed the findings, contributed to the revisions of the manuscript, and approved the final version of the manuscript.
Competing interests
The authors declare no competing interests.
Acknowledgements
This work was supported by Japan Society for the Promotion of Science KAKENHI Grant-in-Aid for Scientific Research (B) Grant Numbers JP25K01454 and JP24K00247, and Watanabe Memorial Foundation for the Advancement of New Technology FY2024 Science and Technology Research Grant Number R6-590. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Supplementary Information
Figure S1: The Pearson correlation between the degree centrality $K_{T,0}$ of technology node $T$ classified by IPC and the average nearest neighbor degree $K_{T,1}$, and the Pearson correlation between the degree centrality $K_{T,0}$ of technology field node $T$ and TCI. In both figures, the black lines indicate the mean values. Each data point is color-coded according to the five classifications defined by Schmoch[Schmoch2008].
Figure S2: Changes in the five-year rolling ranking of the TCI based on 124 three-digit IPC classes for (A) the corporate analysis and (B) the regional analysis. The lines color-coded according to the one-digit IPC classes.
Figure S3: The distribution of weighted patent counts (horizontal axis) for each corporation, with the cumulative distribution function (CDF) of their patents on the vertical axis. The black dashed lines on both axes indicate the excluded region, corresponding to the top 3% of corporations that account for 93% of all patents.